OpenStudy (anonymous):

3x + 9y = 0 x - 9y = 30 A. (-15, -5) B. (5, 11) C. (-14, 5) D. (-1, -15) E. None of the above

OpenStudy (anonymous):

I what to say a

OpenStudy (jordanloveangel):

Simplifying 3x + 9y = -30 Solving 3x + 9y = -30 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-9y' to each side of the equation. 3x + 9y + -9y = -30 + -9y Combine like terms: 9y + -9y = 0 3x + 0 = -30 + -9y 3x = -30 + -9y Divide each side by '3'. x = -10 + -3y Simplifying x = -10 + -3y

OpenStudy (jordanloveangel):

okay to do this first that the 1st eq'n 2x -6y = 30 and graph it. I would graph this line by finding the x and y intercepts x int (let y = 0): (15,0) y int (lets x = 0): (0,-5) next, take the second equation 3x-9y=-27 and graph it as well x int (let y = 0): (-9,0) y int (lets x = 0): (0,3) wherever the lines meet is the sol'n

OpenStudy (anonymous):

So its A

OpenStudy (jordanloveangel):

The solution will be one point on the graph (x,y) where the two lines intercept. The lines should only cross once. The (15,0) and (0,-5) are just two points so you can graph the first line.

OpenStudy (jordanloveangel):

yes

OpenStudy (jordanloveangel):

when these two equations are graphed, they are parallel; which means no solution.

OpenStudy (freckles):

try adding the equations together 3x+9y=0 x- 9y=30 -------

OpenStudy (jordanloveangel):

u get it @$Young_Money$

OpenStudy (freckles):

adding the two equations together simply eliminates the y part and you can solve for x

OpenStudy (anonymous):

Yea

OpenStudy (jordanloveangel):

glad

OpenStudy (jordanloveangel):

if u still need help tag me ight and have a nice day ^_^